# Group Sparse Matrix Optimization for Efficient Quantum State Transformation

### v1

@@ III Problem Formulation
Given an $n$ qubit system, our goal is to find the optimal $n^{2}\times n^{2}$ transformation matrix $\mathbf{X}$ that transforms a given $n^{2}\times 1$ initial state vector $|\psi_{\mathrm{initial}}\rangle$ denoted as $\mathbf{A}$ into the target $n^{2}\times 1$ final state vector $|\psi_{\mathrm{target}}\rangle$ denoted as $\mathbf{C}$. The optimality refers to the extend of sparsity: we intend that $\mathbf{X}$ is as sparse as possible so that quantum operations needed to realize it will be minimized.
We are minimizing the distances between the obtained state $\mathbf{XA}$ and target $\mathbf{C}$, denoted by an objective function $F(\mathbf{X})$. This in turn solves $X|\psi_{\mathrm{initial}}\rangle=|\psi_{\mathrm{target}}\rangle$. Furthermore, one must impose constraints to $\mathbf{X}$. Firstly, $\mathbf{X}$ must be unitary, satisfying $\mathbf{X}^{\dagger}\mathbf{X}=\mathbf{I}$, thus maintaining the probability conservation and reversibility of the quantum operations. Next, to ensure that the target states are properly normalized, for each row or column vector of $\mathbf{X}$, denoted as $\mathbf{x_{r}}$ and $\mathbf{x_{c}}$ respectively, the sum of all its elements equals to one. Note that $r=1\ldots n^{2}$ and $c=1\ldots n^{2}$ represent $i$-th row and $j$-th column of $\mathbf{X}$. For ease of reference, we refer to these constraints as the ‘row/column-wise unit sum constraint’ throughout the paper.
@@
The row/column-wise unit sum constraint can be equivalently expressed as an equality constraint in $F(\mathbf{X})$. In order to enforce row-wise unit sum, multiplication of objective $\mathbf{X}$ with a $n^{2}\times 1$ row $\mathds{1}_{\mathrm{row}}$ should equal $\mathds{1}_{\mathrm{row}}$. Whereas column-wise unit sum can be represented as multiplication of column $1\times n^{2}$ column vector $\mathds{1}_{\mathrm{col}}$ and $\mathbf{X}$ equal $\mathds{1}_{\mathrm{col}}$.
@@
Since objective matrix $\mathbf{X}$ is unitary, i.e. square matrix whose conjugate transpose is equal to its inverse, the rows and columns of $\mathbf{X}$ are orthonormal vectors. Eq. (9) can further be simplified, by merging two equality constraints into one. This property allows us to consider two equivalent constraints: $\mathds{1}\mathbf{X}=\mathds{1}$ and $\mathbf{X}\mathds{1}=\mathds{1}$, where $\mathds{1}$ represent $n^{2}\times n^{2}$ matrix of one. By multiplying $\mathbf{X}$ by $\mathds{1}$ on either side, we preserve the orthonormality of the rows and columns, ensuring that the resulting matrix satisfies the row/column-wise unit sum constraint. Therefore, for the objective function $F(\mathbf{X})$, we employ $\mathds{1}\mathbf{X}=\mathds{1}$ to enforce the unit sum constraint for both rows and columns.
@@ IV ADMM on Quantum State Transformation
In this section we present our method for optimizing group-sparsed unitary matrix for $\mathbf{X}$ in Eq. (12). The incorporation of unitary constraints in matrix optimization is particularly relevant in various quantum problems, including quantum channel estimation Chapeau-Blondeau (2021); Chen et al. (2023), quantum tomography Baldwin et al. (2014); Yang et al. (2017) and quantum state transformation Gleinig and Hoefler (2021); Malvetti et al. (2021); Jaques and Häner (2022); de Veras et al. (2022). Since both unitary constraint and the sparsity term given by either Eq. (6) or Eq. (7) are non-convex, optimizing the objective functions defined in Eq. (11) for general sparsity and Eq. (12) is NP-hard. In this paper, we propose a new approach for solving the problem by introducing two dual variables, denoted as $\mathbf{Z_{1}}$ and $\mathbf{Z_{2}}$ and use ADMM algorithm. These dual variables assist us in effectively managing each constraint in the problem. Additionally, we introduce an auxiliary variable matrix $\mathbf{Y}$ that represents unitary constraint (i.e. $\mathbf{Y}^{\dagger}\mathbf{Y}-\mathbf{I}=\mathbf{0}$). The final objective function to be optimized can be rewritten from Eq. (12) to a new unconstrained objective function:
@@
By minimizing $L$ with respect to $\mathbf{X}$, $\mathbf{Y}$, $\mathbf{Z_{1}}$ and $\mathbf{Z_{2}}$, optimal transformation matrix $\mathbf{X}$ that enhances the minimum values of the objective function while satisfying the sparsity and other structural requirements can be found. Algorithm 1 provides a comprehensive outline for solving Eq. (14).
@@
Algorithm 1 ADMM Algorithm for Quantum State Transformation
@@ IV.1 Updating X Sub-problem
In step 5 of Algorithm 1, the focus is to update $\mathbf{X}$ by seeking its local optimum for each iteration through solving:
@@ IV.2 Updating Y Subproblem
Step 6 of Algorithm 1 aims to solve for $\mathbf{Y}$, which is the variable used to consider the unitary constraint. Since $\mathbf{Y}$ is not involved in sparsity and row/column-wise unit sum constraint, it can be treated as a sub-problem that search for an unitary matrices $\mathbf{Y}$ fulfilling the fidelity condition. Such problem can be considered as the Orthogonal Procrustes problem Schönemann (1966). The Orthogonal Procrustes problem aims to maximize $\langle\mathbf{X},\mathbf{C}\mathbf{A}^{\dagger}\rangle$ which can be solved by Singular Value Decomposition, a matrix factorization routine that breaks down a matrix into three components: a left singular matrix $\mathbf{U}$, a diagonal singular value matrix $\mathbf{S}$, and a right singular matrix $\mathbf{V}$:
@@ V Construction of Quantum Circuit using Optimal Unitary Matrix
      $\begin{array}[]{c|cccccccc}&|000\rangle&|001\rangle&|010\rangle&|011\rangle&|100\rangle&|101\rangle&|110\rangle&|111\rangle\\ \hline\cr|000\rangle&x_{1,1}&x_{1,2}&x_{1,3}&x_{1,4}&x_{1,5}&x_{1,6}&x_{1,7}&x_{1,8}\\ |001\rangle&x_{2,1}&x_{2,2}&x_{2,3}&x_{2,4}&x_{2,5}&x_{2,6}&x_{2,7}&x_{2,8}\\ |010\rangle&x_{3,1}&x_{3,2}&x_{3,3}&x_{3,4}&x_{3,5}&x_{3,6}&x_{3,7}&x_{3,8}\\ |011\rangle&x_{4,1}&x_{4,2}&x_{4,3}&x_{4,4}&x_{4,5}&x_{4,6}&x_{4,7}&x_{4,8}\\ |100\rangle&x_{5,1}&x_{5,2}&x_{5,3}&x_{5,4}&x_{5,5}&x_{5,6}&x_{5,7}&x_{5,8}\\ |101\rangle&x_{6,1}&x_{6,2}&x_{6,3}&x_{6,4}&x_{6,5}&x_{6,6}&x_{6,7}&x_{6,8}\\ |110\rangle&x_{7,1}&x_{7,2}&x_{7,3}&x_{7,4}&x_{7,5}&x_{7,6}&x_{7,7}&x_{7,8}\\ |111\rangle&x_{8,1}&x_{8,2}&x_{8,3}&x_{8,4}&x_{8,5}&x_{8,6}&x_{8,7}&x_{8,8}\\ \end{array}$ (25)
@@ VI Numerical simulation
We generate a collection of complex vectors with size $n^{2}\times 1$ to simulate an entangled state of $n$ qubits. This simulation replicates the process of searching for a transformation matrix that represents a quantum circuit capable of transforming an initial state $\mathbf{A}$ to a target state $\mathbf{C}$, both randomly generated in MATLAB. Additionally, we set the initial transformation matrix $\mathbf{X}_{0}$ as a zero matrix $\mathbf{0}$, recognizing that the initial variables can significantly impact ADMM optimization.
@@ VI.1 Convergence of the Proposed Algorithm
No. of iterations required for convergence $>1000$ $>1000$ $>1000$
Percentage of Non-Zero Elements $<78\%$ $<74\%$ $<60\%$
@@
No. of iterations required for convergence $>100$ $>100$ $>100$ $>100$ $>100$ $>100$
@@ Acknowledgments
This work is supported by the Key-Area Research and Development Program of GuangDong Province (Grant No. 2018B030326001), the National Natural Science Foundation of China (Grant No. 11874312), and the Research Grants Council of Hong Kong (CityU 11304920).
@@ Appendix A Comparison of sparsity for original (unnormalized) vector
No. of iterations required for convergence $>1300$ $>1600$ $>1100$ $>1100$ $>1200$ $>1000$
Percentage of Non-Zero Elements $<76\%$ $<74\%$ $<70\%$ $<65\%$ $<60\%$ $<45\%$
@@
No. of iterations required for convergence $>100$ $>100$ $>200$ $>600$ $>1200$ $>2000$
@@ Appendix B Comparison of sparsity for least square objective problem
No. of iterations required for convergence $>600$ $>900$ $>9500$ $<100$
Percentage of Non-zero Elements $<64\%$ $<27\%$ $<32\%$ $<11\%$
@@
No. of iterations required for convergence $>400$ $>500$ $>3200$ $>4000$
@@ Appendix C Gate Efficiency in Quantum Circuit Decomposition
Average Number of Gates $<26$ $<110$ $<500$ $<2000$ $<8000$

### v2

@@ III Problem Formulation
[…]
We are minimizing the distances between the obtained state $\mathbf{XA}$ and target $\mathbf{C}$, denoted by an objective function $F(\mathbf{X})$. This in turn solves $X|\psi_{\mathrm{initial}}\rangle=|\psi_{\mathrm{target}}\rangle$. Furthermore, one must impose constraints to $\mathbf{X}$. Firstly, $\mathbf{X}$ must be unitary, satisfying $\mathbf{X}^{\dagger}\mathbf{X}=\mathbf{I}$, thus maintaining the probability conservation and reversibility of the quantum operations. Next, to ensure that the target states are properly normalized, for each row or column vector of $\mathbf{X}$, denoted as $\mathbf{x_{r}}$ and $\mathbf{x_{c}}$ respectively, the sum of all its elements equals to one. Note that $r=1\ldots{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2^{n}}$ and $c=1\ldots{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2^{n}}$ represent $i$-th row and $j$-th column of $\mathbf{X}$. For ease of reference, we refer to these constraints as the ‘row/column-wise unit sum constraint’ throughout the paper.
@@
The row/column-wise unit sum constraint can be equivalently expressed as an equality constraint in $F(\mathbf{X})$. In order to enforce row-wise unit sum, multiplication of objective $\mathbf{X}$ with a ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2^{n}}\times 1$ row $\mathds{1}_{\mathrm{row}}$ should equal $\mathds{1}_{\mathrm{row}}$. Whereas column-wise unit sum can be represented as multiplication of column $1\times{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2^{n}}$ column vector $\mathds{1}_{\mathrm{col}}$ and $\mathbf{X}$ equal $\mathds{1}_{\mathrm{col}}$.
@@
[…]
@@ IV ADMM on Quantum State Transformation
In this section we present our method for optimizing group-sparsed unitary matrix for $\mathbf{X}$ in Eq. (12). The incorporation of unitary constraints in matrix optimization is particularly relevant in various quantum problems, including quantum channel estimation Chapeau-Blondeau (2021); Chen et al. (2023), quantum tomography Baldwin et al. (2014); Yang et al. (2017) and quantum state transformation Gleinig and Hoefler (2021); Malvetti et al. (2021); Jaques and Häner (2022); de Veras et al. (2022). Due to the nonconvexity of the unitary constraint, both of the optimization problem defined in Eq. (11) and Eq. (12) are NP-hard. In this paper, we propose a new approach for solving the problem by introducing two dual variables, denoted as $\mathbf{Z_{1}}$ and $\mathbf{Z_{2}}$ and use ADMM algorithm. These dual variables assist us in effectively managing each constraint in the problem. Additionally, we introduce an auxiliary variable matrix $\mathbf{Y}$ that represents unitary constraint (i.e. $\mathbf{Y}^{\dagger}\mathbf{Y}-\mathbf{I}=\mathbf{0}$). The final objective function to be optimized can be rewritten from Eq. (12) to a new unconstrained objective function:
@@
By minimizing $L$ with respect to $\mathbf{X}$, $\mathbf{Y}$, $\mathbf{Z_{1}}$ and $\mathbf{Z_{2}}$, optimal transformation matrix $\mathbf{X}$ that enhances the minimum values of the objective function while satisfying the sparsity and other structural requirements can be found. The ADMM Algorithm for Quantum State Transformation provides a comprehensive outline for solving Eq. (14).
@@
Algorithm ADMM Algorithm for Quantum State Transformation
@@ IV.1 Updating X Sub-problem
In step 5 of our algorithm, the focus is to update $\mathbf{X}$ by seeking its local optimum for each iteration through solving:
@@ IV.2 Updating Y Subproblem
Step 6 of the algorithm aims to solve for $\mathbf{Y}$, which is the variable used to consider the unitary constraint. Since $\mathbf{Y}$ is not involved in sparsity and row/column-wise unit sum constraint, it can be treated as a sub-problem that search for an unitary matrices $\mathbf{Y}$ fulfilling the fidelity condition. Such problem can be considered as the Orthogonal Procrustes problem Schönemann (1966). The Orthogonal Procrustes problem aims to maximize $\langle\mathbf{X},\mathbf{C}\mathbf{A}^{\dagger}\rangle$ which can be solved by Singular Value Decomposition, a matrix factorization routine that breaks down a matrix into three components: a left singular matrix $\mathbf{U}$, a diagonal singular value matrix $\mathbf{S}$, and a right singular matrix $\mathbf{V}$:
@@ V Construction of Quantum Circuit using Optimal Unitary Matrix
[…]
@@ VI Numerical simulation
We generate a collection of complex vectors with size ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}2^{n}}\times 1$ to simulate an entangled state of $n$ qubits. This simulation replicates the process of searching for a transformation matrix that represents a quantum circuit capable of transforming an initial state $\mathbf{A}$ to a target state $\mathbf{C}$, both randomly generated in MATLAB. Additionally, we set the initial transformation matrix $\mathbf{X}_{0}$ as a zero matrix $\mathbf{0}$, recognizing that the initial variables can significantly impact ADMM optimization.
@@ VI.1 Convergence of the Proposed Algorithm
No. of iterations required for convergence >1000>1000 >1000>1000 >1000>1000
Percentage of Non-Zero Elements $78\% <$$74\% <$$60\%$
@@
No. of iterations required for convergence >100>100 >100>100 >100>100 >100>100 >100>100 >100>100
@@ Acknowledgments
This work is supported by the National Natural Science Foundation of China (Grant No. 11874312), the Research Grants Council of Hong Kong (CityU 11304920), the Guangdong Provincial Quantum Science Strategic Initiative (Grant No. GDZX2203001, GDZX2303007), and the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0302300).
@@ Appendix A Comparison of sparsity for original (unnormalized) vector
No. of iterations required for convergence >1300>1300 >1600>1600 >1100>1100 >1100>1100 >1200>1200 >1000>1000
Percentage of Non-Zero Elements $76\% <$$74\% <$$70\% <$$65\% <$$60\% <$$45\%$
@@
No. of iterations required for convergence >100>100 >100>100 >200>200 >600>600 >1200>1200 >2000>2000
@@ Appendix B Comparison of sparsity for least square objective problem
No. of iterations required for convergence >600>600 >900>900 >9500>9500 <100<100
Percentage of Non-zero Elements $64\% <$$27\% <$$32\% <$$11\%$
@@
No. of iterations required for convergence >400>400 >500>500 >3200>3200 >4000>4000
@@ Appendix C Gate Efficiency in Quantum Circuit Decomposition
Average Number of Gates <26<26 <110<110 <500<500 <2000<2000 <8000<8000